Solve by completing the square
x^2=5x+2
Any help is appreciated

Answer by stanbon(75887)

Solve by completing the square
x^2=5x+2
------------
x^3-5x = 2
Complete the square on the x-terms and balance the equation:
x^2-5x+(5/2)^2 = 2 + (5/2)^2
Factor the left side and simplify the right side:
(x-(5/2))^2 = 33/4
Take the square root of both sides:
(x-(5/2)) = (sqrt33)/2 or (x-(5/2)) = (-sqrt33)/2
Solve for x:
x = [5+sqrt33]/2 or x = [5-sqrt33]/2
=================
Cheers,
Stan H.